is 3 a solution to the equation 2x-4=x-1
Found 2 solutions by RAY100, Alan3354:
Answer by RAY100(1637)

2x-4 = x-1,,,,,,,add 4 to both sides
.
2x-4+4 = x-1+4
.
2x=x+3,,,,,,now, substract x both sides
.
2x-x = x+3-x
.
x=3,,,,,answer
.
.
check
.
2x-4 = x-1,,,,subst x=3
.
2(3)-4 = (3)-1
.
2=2,,,,,,ok
,

Answer by Alan3354(69443)

Yes, it's the only solution.